### v3

# Think-to-Talk or Talk-to-Think? When LLMs Come Up with an Answer in Multi-Hop Arithmetic Reasoning

Abstract
This study investigates the incremental, internal problem-solving process of language models (LMs) with arithmetic multi-hop reasoning as a case study. We specifically investigate when LMs internally resolve sub/whole problems through first reading the problem statements, generating reasoning chains, and achieving the final answer to mechanistically interpret LMs’ multi-hop problem-solving process. Our experiments reveal a systematic incremental reasoning strategy underlying LMs. They have not derived an answer at the moment they first read the problem; instead, they obtain (sub)answers while generating the reasoning chain. Therefore, the generated reasoning chains can be regarded as faithful reflections of the model’s internal computation.
[…]
100
Think-to-Talk or Talk-to-Think?
When LLMs Come Up with an Answer in Multi-Hop Arithmetic Reasoning
Keito Kudo1,2, Yoichi Aoki1,2, Tatsuki Kuribayashi3, Shusaku Sone1 Masaya Taniguchi2, Ana Brassard2,1, Keisuke Sakaguchi1,2, Kentaro Inui3,1,2 1Tohoku University, 2RIKEN, 3MBZUAI {keito.kudo.q4, youichi.aoki.p2, sone.shusaku.r8}@dc.tohoku.ac.jp, {tatsuki.kuribayashi, kentaro.inui}@mbzuai.ac.ae, keisuke.sakaguchi@tohoku.ac.jp, {masaya.taniguchi, ana.brassard} @riken.jp
1 Introduction
An explanation may be produced in two modes: as a post hoc explanation to a predetermined conclusion (Think-to-Talk) or by the process of reaching a conclusion while explaining (Talk-to-Think). An analogy applies to large language models (LLMs) using chain-of-thought (CoT; Wu et al., 2023a) reasoning: is generated CoT reasoning chain a post-hoc explanation, or does it reflect real-time step-by-step solving? This question is particularly relevant to the (mechanistic) interpretablity of LLMs — how models incrementally resolve the multi-hop problem, in other words, how information internally flows to reach the final decision.
In this study, we identify such internal reasoning patterns of LLMs. There may be multiple possible internal strategies for LLMs under the CoT scenario, i.e., first feed the problem statements and then have models generate full reasoning chains. One presumably hard strategy for LLMs would be to reach the final answer even during the first pass of problem statements before CoT generation, and then while generating CoT reasoning chains, the model simply refers to their predetermined answers as a post-hoc explanation (think-to-talk mode). The opposite will be that models do nothing during the first pass of the problem statements and start solving the problem after CoT generation begins (talk-to-think mode). Such different internal mechanisms can not be distinguished by just observing the model outputs; rather, one has to interpret and intervene in their internal representations during their multi-hop CoT-style reasoning.
Figure 1: Using linear probes, we investigated at which time during the LLM’s problem-solving process it is possible to determine the values of each variable, illustrating the model’s problem-solving process. Our analysis indicates that LMs come up with (sub)answers during CoT (second pass). This conclusion is also consistent with the findings from the causal experiments in §˜5.
In our first experiments, we apply linear probes to model internals at each layer at each timestep to determine when answers are reached internally. We prepared controlled testbeds of symbolic arithmetic reasoning tasks and observed whether trained probes could accurately predict and control the values of specific variables (Figure 1). By comparing accuracies across each timestep, one can observe at which point models internally start being informative to the probes, illustrating the model’s internal reasoning flow.
The results reveal common patterns across models. They have not derived an answer at the moment they first read the problem (first-pass); instead, they obtain (sub)answers while generating the reasoning chain. These tendencies are consistent and systematic across our different testbeds.
Based on the above finding, we further conducted causal intervention analyses to clarify the causal relationship between the model’s internal representations and the final answer (§˜5). We found that, when generating the reasoning chain, the model’s internal problem-solving process exhibits a recency bias, relying heavily on the information in the immediately preceding portion of the chain. In other words, the generated reasoning chains can be regarded as faithful reflections of the model’s internal computation.  [^1] Code and data will be made available upon acceptance.
Level Input Output #Step #Stack #Dist.
1 ${\underline{\mathrm{A}=1+\mathrm{B}}}_{\hskip 1.42271pt\mathbf{-3}}$, ${\underline{\mathrm{B}=2}}_{\hskip 1.42271pt\mathbf{-2}}$; ${\underline{\mathrm{A}=?}}_{\hskip 1.42271pt\mathbf{-1}}$ ${\underline{\mathrm{A}=1+\mathrm{B}}}_{\hskip 1.42271pt\mathbf{0}}$, ${\underline{\mathrm{B}=2}}_{\hskip 1.42271pt\mathbf{1}}$, ${\underline{\mathrm{A}=1+\mathrm{B}}}_{\hskip 1.42271pt\mathbf{2}}$, ${\underline{\mathrm{A}=1+2}}_{\hskip 1.42271pt\mathbf{3}}$, ${\underline{\mathrm{A}=3}}_{\hskip 1.42271pt\mathbf{4}}$ 1 1 0
2 ${\underline{\mathrm{A}=2+3}}_{\hskip 1.42271pt\mathbf{-3}}$, ${\underline{\mathrm{B}=1+\mathrm{A}}}_{\hskip 1.42271pt\mathbf{-2}}$; ${\underline{\mathrm{B}=?}}_{\hskip 1.42271pt\mathbf{-1}}$ ${\underline{\mathrm{B}=1+\mathrm{A}}}_{\hskip 1.42271pt\mathbf{0}}$, ${\underline{\mathrm{A}=2+3}}_{\hskip 1.42271pt\mathbf{1}}$, ${\underline{\mathrm{A}=5}}_{\hskip 1.42271pt\mathbf{2}}$, ${\underline{\mathrm{B}=1+\mathrm{A}}}_{\hskip 1.42271pt\mathbf{3}}$, ${\underline{\mathrm{B}=1+5}}_{\hskip 1.42271pt\mathbf{4}}$, ${\underline{\mathrm{B}=6}}_{\hskip 1.42271pt\mathbf{5}}$ 2 0 0
3 ${\underline{\mathrm{A}=1+\mathrm{B}}}_{\hskip 1.42271pt\mathbf{-3}}$, ${\underline{\mathrm{B}=2+3}}_{\hskip 1.42271pt\mathbf{-2}}$; ${\underline{\mathrm{A}=?}}_{\hskip 1.42271pt\mathbf{-1}}$ ${\underline{\mathrm{A}=1+\mathrm{B}}}_{\hskip 1.42271pt\mathbf{0}}$, ${\underline{\mathrm{B}=2+3}}_{\hskip 1.42271pt\mathbf{1}}$, ${\underline{\mathrm{B}=5}}_{\hskip 1.42271pt\mathbf{2}}$, ${\underline{\mathrm{A}=1+\mathrm{B}}}_{\hskip 1.42271pt\mathbf{3}}$, ${\underline{\mathrm{A}=1+5}}_{\hskip 1.42271pt\mathbf{4}}$, ${\underline{\mathrm{A}=6}}_{\hskip 1.42271pt\mathbf{5}}$ 2 1 0
4 ${\underline{\mathrm{A}=1+\mathrm{B}}}_{\hskip 1.42271pt\mathbf{-4}}$, ${\underline{\mathrm{B}=2+3}}_{\hskip 1.42271pt\mathbf{-3}}$, ${\underline{\mathrm{C}=4+5}}_{\hskip 1.42271pt\mathbf{-2}}$; ${\underline{\mathrm{A}=?}}_{\hskip 1.42271pt\mathbf{-1}}$ ${\underline{\mathrm{A}=1+\mathrm{B}}}_{\hskip 1.42271pt\mathbf{0}}$, ${\underline{\mathrm{B}=2+3}}_{\hskip 1.42271pt\mathbf{1}}$, ${\underline{\mathrm{B}=5}}_{\hskip 1.42271pt\mathbf{2}}$, ${\underline{\mathrm{A}=1+\mathrm{B}}}_{\hskip 1.42271pt\mathbf{3}}$, ${\underline{\mathrm{A}=1+5}}_{\hskip 1.42271pt\mathbf{4}}$, ${\underline{\mathrm{A}=6}}_{\hskip 1.42271pt\mathbf{5}}$ 2 1 1
5 ${\underline{\mathrm{A}=1+\mathrm{B}}}_{\hskip 1.42271pt\mathbf{-4}}$, ${\underline{\mathrm{B}=2+\mathrm{C}}}_{\hskip 1.42271pt\mathbf{-3}}$, ${\underline{\mathrm{C}=1+2}}_{\hskip 1.42271pt\mathbf{-2}}$; ${\underline{\mathrm{A}=?}}_{\hskip 1.42271pt\mathbf{-1}}$ ${\underline{\mathrm{A}=1+\mathrm{B}}}_{\hskip 1.42271pt\mathbf{0}}$, ${\underline{\mathrm{B}=2+\mathrm{C}}}_{\hskip 1.42271pt\mathbf{1}}$, ${\underline{\mathrm{C}=1+2}}_{\hskip 1.42271pt\mathbf{2}}$, ${\underline{\mathrm{C}=3}}_{\hskip 1.42271pt\mathbf{3}}$, ${\underline{\mathrm{B}=2+\mathrm{C}}}_{\hskip 1.42271pt\mathbf{4}}$, ${\underline{\mathrm{B}=2+3}}_{\hskip 1.42271pt\mathbf{5}}$, ${\underline{\mathrm{B}=5}}_{\hskip 1.42271pt\mathbf{6}}$, ${\underline{\mathrm{A}=1+\mathrm{B}}}_{\hskip 1.42271pt\mathbf{7}}$, ${\underline{\mathrm{A}=1+5}}_{\hskip 1.42271pt\mathbf{8}}$, ${\underline{\mathrm{A}=6}}_{\hskip 1.42271pt\mathbf{9}}$ 3 2 0
Table 1: Examples of arithmetic reasoning tasks used in our experiments at each complexity level. #Step indicates the number of required operations to reach the final answer. #Stack indicates how many variables’ values are not immediately determined in their first appearing equation. #Dist. is the number of unnecessary distractor equations. The number (e.g., $\mathbf{-3}$) indicated in the lower right corner of each equation represents the equation’s position. This position is used as a reference point for calculating $t^{*}_{\mathrm{eq}}$ in §˜4.2.
2 Related Work
Multi-stage human incremental language comprehension.
Humans sometimes make several attempts with different strategies or mindsets, particularly when resolving a complex task. A common view may be, for example, that humans first adopt a shallow, fast solution, and once it fails, switch to a more expensive, presumably accurate one. Such a multi-stage processing is even related to the recent debate on the cognitive plausibility of LLMs in computational psycholinguistics Oh and Schuler (2023); Shain et al. (2024); Kuribayashi et al. (2024); Gruteke Klein et al. (2024). LLMs have now been criticized because they can not estimate human cognitive costs incurred, possibly by the multi-stage nature of human sentence processing (i.e., reanalysis) van Schijndel and Linzen (2021); Huang et al. (2024); humans re-read the sentence again from earlier points as an additional trial when facing difficulties in comprehension, but LLMs do not explicitly have such a multi-stage mechanism. Although this study is not focused on sentence processing or LMs’ humanlikeness, one critical question in this line is whether and how LLMs switch their reasoning strategy for the same problem through multiple trials, and how to track their dynamic internal states. In our case, we simplify the setting to forced-decode the common chain-of-thought style format as two-time trials of the problem, where an LM first reads the problem statement and then (re-)analyzes the problem while generating the reasoning chain as well as copying the problem statement again. Here, we analyze what kind of process LLMs perform, particularly in the first pass of the problem statement, and how their initial processing is related to their later generation of reasoning chain and answer. Our results suggest some surprising behaviors; models systematically come up with answers to simple subproblems in the first pass, but these computations are not reused in the second pass of CoT-style reasoning, suggesting some redundancy in their internal reasoning.
Interpreting multi-hop reasoning in language models.
Interpreting internal mechanisms of LMs has been actively investigated Conneau et al. (2018); Tenney et al. (2019); Niven and Kao (2019); nostalgebraist (2020); Geva et al. (2023); Lieberum et al. (2024); Ghandeharioun et al. (2024); Ferrando and Voita (2024). They revealed, for example, specialized attention heads responsible for specific operations (Cabannes et al., 2024) or decision-making, copying, and induction (Dutta et al., 2024). With a more concrete example, Yang et al. (2024c) showed that, even during the first pass of the roblem statements such as The mother of the singer of Thriller is ___, language models first resolve a bridge entity, Stevie Wonder in this case, then identify the final answer. This study is more focused on the difference between the first pass of the problem statements (before CoT generation) and their second pass involving explicit problem solving (while CoT generation).
Arithmetic representations in LLMs.
How models handle numerical information has also been closely studied. For instance, Heinzerling and Inui (2024) used partial least squares regression Wold et al. (2001) to demonstrate that numeric attributes, such as birth years and population numbers, are encoded as monotonic representations in the activation space of LLMs and can be manipulated with interventions. In turn, Stolfo et al. (2023) showed that, in autoregressive LMs, the operations and numerical information necessary for solving quantitative reasoning are processed in the lower layers of the model, and these results are used by the attention layers to predict the final calculation outcomes. Zhu et al. (2025) studied the representation of numbers in language models’ hidden states during single-hop arithmetic tasks (e.g., What is the sum of 12 and 34?). Their analysis revealed that numerical information is encoded linearly within the hidden states and demonstrated that individual digits could be manipulated independently. In this study, we add to this literature by introducing incremental arithmetic problem solving, i.e., what numerical information is contained in the model’s hidden states at each time step of multi-hop arithmetic reasoning.
Model interpretability methods.
Linear probing Alain and Bengio (2017b) is one of the representative methods for analyzing the internal representations of neural models—a small model predicts a specific feature from them, thereby determining whether the input contains information about that feature. In this study, we use them to derive the models’ intermediate answers. The causality with the model’s output can be further verified by examining if a model’s predictions change when the hidden states are intervened Li et al. (2023); Wu et al. (2023b). One representative intervention method is activation patching Vig et al. (2020); Meng et al. (2022); Zhang and Nanda (2024), where hidden states obtained from one model instance are transplanted onto another during inference to change its predictions. Such techniques can be applied as a way to control model behavior in practical scenarios such as mitigating inherent biases Zhao et al. (2019); Ganguli et al. (2023); Yang et al. (2024b). Here, we employ activation patching to validate the plausibility of the probing results.
3 General settings
3.1 Arithmetic problems
We prepared a dataset of multi-hop arithmetic problems similarly to Kudo et al. (2023) and Yu (2025). Each sample is a string of assignments (e.g., A=1) and operations (e.g., B=1+3 or B=1+A) ending with a query for a variable’s value (e.g., B=?). We also defined five complexity levels, depending on (i) how many equations need to be resolved to reach the answer (#Step in Table 1), (ii) how many variables’ values cannot be immediately resolved (and thus pended to a stack) in their first appearance when incrementally reading the problem from left to right (#Stack), (iii) and the number of unnecessary distractor equations (#Dist.). For example, in the Level 5 example in Table 1, where #Step is three and #Stack is two, calculating A requires at least three steps of reasoning: C(=1+2)=3, B(=2+3)=5, and then A(=1+5)=6, and two variables need to be resolved before reaching A: B and C.
Notation.
Formally, let $v$ denote a variable name sampled from the 26 letters of the English alphabet $\Sigma=\{\text{a},\text{b},\text{c},\dots,\text{z}\}$, and $d$ a number sampled from the set of decimal digits $D=\{0,1,2,\ldots,9\}$. Each instance consists of multiple equations $e_{1},e_{2},\cdots,e_{n}]$ followed by a final query $q$. Each equation follows the format $v=d$, $v=d\pm d$, or $v=d\pm v$. We denote $i$-th variable to appear within an instance from the left as $v_{i}$. E.g., in the Level 5 example in Table 1, $v_{1}=\texttt{A}$, $v_{2}=\texttt{B}$, and $v_{3}=\texttt{C}$. [^2] For ease of reading, all examples throughout this paper use the uppercased variables A, B, C, and only the operator +. The value assigned to a variable $v_{i}$ is denoted as $\{v_{i}\}\in D$.
Generation rules.
We ensure that $\{v_{i}\}$ for any $v_{i}$ is also a single-digit number, and $\{v_{i}\}$ is constant within the same instance (i.e., we exclude cases such as A=1+2,A=B+2,B=6). All samples of the same complexity level follow the exact same format except for the actual numbers, variable names, and operators, meaning that $\{v_{i}\}$ in each level can be obtained with exactly the same abstract procedure. E.g., first calculate $\{v_{3}\}$, and then calculate $\{v_{2}\}$ with $\{v_{3}\}$, and then finally calculate $\{v_{1}\}$ with $\{v_{2}\}$. Non-duplicated instances are created for each level by varying the variable names, numbers, and operators appearing in the equations. We created 12,000 unique instances in total, of which 10,000 are used for training probing classifiers and 2,000 for testing their accuracy. To prevent the probe itself from performing arithmetic, we constructed the training and test sets so that no duplicate arithmetic expressions appear. [^3] For example, if 1+2 appears in the training set, then 1+2 does not appear in the test set.
3.2 CoT-style inference
Henceforth, we refer to the part before CoT (explained in 3.1) as the Input and the CoT-reasoning part as the Output of an instance. The Output, as shown in the right part of Table 1, is also a sequence of equations in the same format as the Input and can be split into intermediate steps $z$ and a final answer $y$. For example, for the Level 1 instance in Table 1 (topmost), x=x=A=1+B, B=2,” z=z=A=1+B, B=2, A=1+2, A=,” and y=y=3.” That is, the task is to generate intermediate steps $z$ and derive a final answer $y$ given an Input $x$ and a demonstration of three examples. As a sanity check, we confirmed that the target models could follow the expected Output format and solve the tasks with nearly 100% accuracy in this setting (see §˜A.1 for more details).
We denote a token position within the entire concatenated sequence $x\oplus z\oplus y$ with $\in\mathbb{Z}. tt$ is relative to CoT; that is, $t$ is zero where CoT begins, negative within the Input, and positive within the Output. Similarly, we assign an equation position $t_{\mathrm{eq}}\in\mathbb{Z}$ to each equation in the Input and Output (subscripts on the underlines in Table 1).
4 Probing
When do models solve (sub)problems in CoT-style reasoning? Do they (i) finalize reasoning during the Input stage, with the CoT as a post-hoc explanation (Think-to-Talk), or (ii) solve the task step-by-step during CoT generation (Talk-to-Think)? We address this by examining where the final answer, or the necessary information for it, emerges in the model’s internal representations using linear probes.
4.1 Training settings for linear probes
We train a linear probe Alain and Bengio (2017a) for an $l$-layer LLM ($l=0$ for the input embedding layer). To identify where the LLM solved a particular (sub)problem, we train a separate probe for each combination of token position $t\in\mathbb{Z}$, layer depth $l\in\mathbb{N}$, and $v_{i}\in\Sigma$ in each level of the problem. Specifically, given a model’s $d$-dimentional hidden state $\bm{h}_{t,l}\in\mathbb{R}^{d}$, the probing classifier $f_{t,l,v_{i}}(\cdot):\mathbb{R}^{d}\to D$ predicts $\{v_{i}\}$. That is, for each $(t,l)$, we first obtained 10,000 of $\bm{h}_{t,l}$ from training instances and then evaluated the accuracy of $f_{t,l,v_{i}}(\cdot)$ for each $v_{i}$ with 2,000 hidden states from test instances and the correct $\{v_{i}\}$. If a probe $f_{t,l,v_{i}}(\cdot)$ achieves high accuracy, this suggests that $\{v_{i}\}$ is already computed at the corresponding position ($t$ and $l$). Figure 2 illustrates the probing results with a Level 3 task, where, for example, the value of B can be extracted within Input and thus already computed before CoT begins.
The probe $f_{t,l,v_{i}}(\cdot)$ is a single linear transformation; that is, the probe is applied to $\bm{h}_{t,l}$ as follows:
$\hat{\$\{v_{i}\}}_{t,l,}$ $=f_{t,l,v_{i}}(\bm{h}_{t,l})$ (1)
$=\operatorname*{arg\,max}_{D}\bm{W}_{t,l,v_{i}}\bm{h}_{t,l}+\bm{b}_{t,l,v_{i}}\;\;\mathrm{,}$
where, $\bm{W}_{t,l,v_{i}}\in\mathbb{R}^{|D|\times d}$ and $\bm{b}_{t,l,v_{i}}\in\mathbb{R}^{|D|}$ are the weight and bias parameters of the probe, respectively. The symbol $\hat{\cdot}$ is used to refer to the model’s estimate. We train the probes using stochastic gradient descent Robbins (1951) to minimize the cross entropy loss. The hyperparameters are listed in Table 5 in the appendix.
Figure 2: Probing results for Qwen2.5-7B at the task Level 3. The heatmaps in the lower section represent the accuracy of probes computed on the evaluation set. Each cell shows the probing accuracies in each token $t$, layer $l$. The upper part indicates the maximum probing accuracy achieved at each token position $t$. The input sequence below the line graphs is just an example; in the actual evaluation set, each variable name, number, and operator are randomly sampled from $(D,\Sigma,\{+,-\})$.
4.2 Evaluation metrics
The probing results from all the token positions $t$ and layers $l$ are aggregated as follows:
$t^{*}(v_{i})$ $=\min\{t\mid\max_{l}\mathrm{acc}(t,l,v_{i})>\tau\}\;\;\mathrm{,}$ (2)
where $\mathrm{acc}(t,l,v_{i})\in[0,1]$ indicates the accuracy of a probing classifier $f_{t,l,v_{i}}$. The $t^{*}(v_{i})\in\mathbb{Z}$ indicates when was the first time the probing classifier achieved a reasonable accuracy above $\tau$ (= 0.90 in our study [^4] See Appendix A.2 for results at different thresholds.) for the variable $v_{i}$. As a more coarse but comprehensive value, we also report $t_{\mathrm{eq}}^{*}(v_{i})$ indicating which equation $t_{\mathrm{eq}}$ the $t^{*}(v_{i})$ falls into. Given that the $t$ (and $t_{\mathrm{eq}}$) is relative to the CoT-beginning position, if $t_{\mathrm{(eq)}}^{*}(v_{i})$ is negative, the value $\{v_{i}\}$ is computed before CoT begins. This is the case for $t^{*}(i.e.,\mathrm{\texttt{B}})$ in Figure 2, based on the spike around $t_{\mathrm{eq}}$=-2 in the upper line graph.
We also report two types of accuracy:
$\mathrm{Acc}_{\mathrm{\prec CoT}}(v_{i})$ $=\max_{t<0,l}\mathrm{acc}(t,l,v_{i})\text{,}$ (3)
$\mathrm{Acc}_{\mathrm{\succ CoT}}(v_{i})$ $=\max_{t\geq 0,l}\mathrm{acc}(t,l,v_{i})\text{.}$ (4)
If $\mathrm{Acc}_{\mathrm{\prec CoT}}(v_{i})$ is sufficiently high, $\{v_{i}\}$ is resolved internally before CoT begins (Think-to-Talk mode). Conversely, if $\mathrm{Acc}_{\mathrm{\prec CoT}}(v_{i})$ is low and $\mathrm{Acc}_{\mathrm{\succ CoT}}(v_{i})$ is high, the answer is derived while performing CoT reasoning (Talk-to-Think mode).
4.3 Experimental results
Across task complexity levels.
We first analyze Qwen2.5-7B Qwen Team (2024) across the five task levels. Table 2 shows $t_{\mathrm{eq}}^{*}$ for each variable as well as the lower bounds of $t^{\dagger}_{\mathrm{eq}}$, which can be computed with a greedy resolver of equations. In most cases, regardless of #Steps or task level, $t^{*}_{\mathrm{eq}}>0$. The exceptions are $v_{2}$ in level 1 and $v_{3}$ in level 4. $v_{2}$ in level 1 has #Step =0=0 and is a variable whose value requires no computation to derive. $v_{3}$ in level 4 is a distractor and is not required to derive the final answer. Therefore, we find that the variables that are required to derive the final answer and require computation are all solved after CoT begins. In summary, we find that the model solved all subproblems necessary to derive the final answer during CoT, and that the Talk-to-Think mode is dominant.
Variable When ($\downarrow$) $\mathrm{Acc.}$ ($\uparrow$)
Level variable #Step $t_{\mathrm{eq}}^{*}$ $t_{\mathrm{eq}}^{\dagger}$ $\mathrm{\prec CoT}$ $\mathrm{\succ CoT}$
1 $v_{1}$ ($\mathrm{A}$) 1 4 -2 0.36 1
$v_{2}$ ($\mathrm{B}$) 0 -2 -2 1 1
2 $v_{1}$ ($\mathrm{A}$) 1 2 -3 0.49 1
$v_{2}$ ($\mathrm{B}$) 2 5 -2 0.21 1
3 $v_{1}$ ($\mathrm{A}$) 2 5 -2 0.18 1
$v_{2}$ ($\mathrm{B}$) 1 2 -2 0.50 1
4 $v_{1}$ ($\mathrm{A}$) 2 5 -3 0.17 1
$v_{2}$ ($\mathrm{B}$) 1 2 -3 0.48 1
$v_{3}$ ($\mathrm{C}$) 1 N/A -2 0.44 0.24
5 $v_{1}$ ($\mathrm{A}$) 3 9 -2 0.18 1
$v_{2}$ ($\mathrm{B}$) 2 6 -2 0.23 1
$v_{3}$ ($\mathrm{C}$) 1 3 -2 0.51 1
Table 2: The results of Qwen2.5-7B on the five levels. The $t^{*}_{\mathrm{eq}}$ is the time when the model comes up with the correct answer (see §˜4.2). The $t_{\mathrm{eq}}^{\dagger}$ column indicates the lower bound of $t^{*}_{\mathrm{eq}}$ score. The $\mathrm{\prec CoT}$ and $\mathrm{\succ CoT}$ scores correspond to the accuracies introduced in §˜4.2. N/A indicates that the threshold $\tau$ was not exceeded at any position $t$.
Across models.
We further analyzed ten models listed in Table 3 on the Level 3 task. Same results as Qwen2.5-7B are generally obtained across various models, enhancing the generality of our obtained findings. These results are consistently observed across other tasks and irrespective of the threshold $\tau$. For more details, see §˜A.2.
When ($\downarrow$) $\mathrm{Acc}$ ($\uparrow$)
Variable $^{*}_{\mathrm{eq}} tt^{*}$ $\mathrm{\prec CoT}$ $\mathrm{\succ CoT}$
Qwen2.5 (7B) $v_{1}$ ($\mathrm{A}$) 5 36 17.9 100
Qwen Team (2024) $v_{2}$ ($\mathrm{B}$) 2 16 50.5 100
Qwen2.5 (14B) $v_{1}$ ($\mathrm{A}$) 5 35 17.8 100
Qwen Team (2024) $v_{2}$ ($\mathrm{B}$) 2 16 50.5 100
Qwen2.5 (32B) $v_{1}$ ($\mathrm{A}$) 5 36 17.8 100
Qwen Team (2024) $v_{2}$ ($\mathrm{B}$) 2 15 67.4 100
Qwen2.5-Math (7B) $v_{1}$ ($\mathrm{A}$) 5 35 18.6 100
Yang et al. (2024a) $v_{2}$ ($\mathrm{B}$) 2 15 56.1 100
Yi1.5 (9B) $v_{1}$ ($\mathrm{A}$) 5 41 17.8 100
Young et al. (2024) $v_{2}$ ($\mathrm{B}$) 2 18 36.9 100
Yi1.5 (34B) $v_{1}$ ($\mathrm{A}$) 5 41 22.4 100
Young et al. (2024) $v_{2}$ ($\mathrm{B}$) 2 18 37.4 100
Llama3.1 (8B) $v_{1}$ ($\mathrm{A}$) 5 35 26.0 100
Dubey et al. (2024) $v_{2}$ ($\mathrm{B}$) 2 16 29.6 100
Llama3.2 (3B) $v_{1}$ ($\mathrm{A}$) 5 36 17.8 93.2
Dubey et al. (2024) $v_{2}$ ($\mathrm{B}$) 2 17 33.2 95.4
Mistral-Nemo (12B) $v_{1}$ ($\mathrm{A}$) 5 36 17.8 100
Mistral AI Team (2024) $v_{2}$ ($\mathrm{B}$) 2 16 28.9 100
Table 3: Results for various models on the task Level 3. The $t^{*}$ column shows the token-wise time (described in §˜4.2), and the other columns are the same as Table 2. The $t^{*}$ and $t^{*}_{\mathrm{eq}}$ scores that are the same as their lower bounds are bolded.
Figure 3: Overview of the causal intervention experiment. First, we perform normal inference (Clean run) and cache its hidden states. Subsequently, we evaluate whether the output changes by replacing some of the hidden states of a model solving a different problem with the cached hidden states.
Figure 4: Success rates for each grid when the final answer $y$ $({\underline{{\color[rgb]{.5,.5,.5}\mathrm{A}=}\textbf{6}}}_{\hskip 1.42271pt\mathbf{5}}$) is the target token.
Figure 5: The upper part is the accuracy of probs, as shown in Figure 2. The lower part is the result of max pooling the Success rates from Figure 4 in the layer direction.
4.4 Analysis
Distractors.
In task Level 4 (see Table 1), $v_{3}$ (C) is a distractor, that is, $\{v_{3}\}$ is not necessary to derive the final answer. The models can infer this fact from the in context examples. According to Table 2, the $\mathrm{Acc}(v_{3})$ in Level 4 was at most 44%, a relatively low accuracy. From this result, we can see that, unlike the variables required to derive the final answer, $v_{3}$ is not encoded in a simple form that can be extracted by a linear transformation alone. This suggests the possibility that the model employs an efficient internal mechanism that does not derive variables unnecessary for obtaining the final answer. It is also consistent with the finding that a Talk-to-Think mode, in which computation is performed during CoT, is dominant.
5 Causal interventions
Motivation.
From our probing experiments in §˜4, we found that the sub-problems needed to produce the final answer are resolved after CoT-style generation begins. However, how this information propagates through the model remains unclear. To shed light on this flow of information, we intervene in the hidden states of the model during reasoning and analyze how these interventions affect the model’s outputs.
5.1 Settings
Activation patching.
We employ activation patching Vig et al. (2020); Meng et al. (2022); Zhang and Nanda (2024), which is a widely adopted technique for causal intervention analysis. To inspect the causal relationship between specific hidden states $\bm{h}_{t,l}$ and a final answer $\hat{y}$, we compare two generation scenarios: (i) the ordinary inference and (ii) the intervened inference. In the latter scenario, we replace the specific hidden states $\bm{h}_{t,l}$ with other variants $\tilde{\bm{h}}_{t,l}$ obtained from the same model but with a different input $\tilde{x}$ (Clean run in Figure 3).
The input $x$ and $\tilde{x}$ have different correct answer $y$ and $\tilde{y}$ as well as different chains $z$ and $\tilde{z}$, respectively. For example, for the triple (x=x=``A=1+B,B=2+4;A=?,” z=z=``A=1+B,B=2+4,B=6,A=1+B,A=1+6,A=7,” y=y=7), one may use ($\tilde{x}$=``A=2+B,B=1+3;A=?,” $\tilde{z}$=``A=2+B,B=1+3,B=4,A=2+B,A=2+4,A=6,” $\tilde{y}$=6). If the model’s output turns from $y$ into $\tilde{y}$ or $z_{t}$ into $\tilde{z_{t}}$ due to the intervention to $\bm{h}_{t,l}$ with $\tilde{\bm{h}}_{t,l}$, we can confirm the causal relationship between $\bm{h}_{t,l}$ and the original answer $y$. We denote the model’s final output without intervention as $\hat{\tilde{y}}$ and that with intervention as $\hat{y}$, respectively ($\tilde{y}$ and $y$ denote respective gold answers). In the same way, we denote the generated reasoning chain without intervention as $\hat{\tilde{z_{t}}}$ and that with intervention as $\hat{z_{t}}$, respectively.
Evaluation metrics.
We report Success Rate as a metric for this experiment. The Success rate indicates how frequently (%) the intervened output $\hat{\tilde{y}}$ aligns with the correct answer $\tilde{y}$. For reasoning chains, we report the Success rate for $\hat{\tilde{z_{t}}}$ as well.
Patching targets.
We specifically focus on Level 3 tasks and Qwen2.5-7B. Inspired by sliding window patching Zhang and Nanda (2024), we partition the hidden states into coarse grids, corresponding to each equation and every four layers, and perform activation patching on each grid separately (illustrated in Figure 5). [^5] All the hidden states in each grid are intervened at once. For every grid, we compute the Success rate by applying activation patching. We also examine multiple target tokens, specifically, at (i) the end of the equation 2 ($z_{17}$ in ${\underline{{\color[rgb]{.5,.5,.5}\mathrm{B}=}\textbf{5}}}_{\hskip 1.42271pt\mathbf{2}}$ in Figure 2), (ii) the end of the equation 4 ($z_{32}$ in ${\underline{{\color[rgb]{.5,.5,.5}\mathrm{A}=1+}\textbf{5}}}_{\hskip 1.42271pt\mathbf{4}}$ in Figure 2), and (iii) the final answer ($y$). When we apply activation patching, we generate the only target token with greedy decoding while forced-decoding the context. Note that the above equations are examples. We create a test set of 2,000 instances for evaluations.
5.2 Results.
Figure 4 shows the Success rate for each grid when the final answer $y$ is the target token. We also show the results for the target tokens $z_{17}$ and $z_{32}$ in Appendix C. Figure 5 summarizes the max success rate among layers for each target token, and probing accuracy (same as the line graph in Figure 2).
[…]
6 Conclusions
We conducted causal probing analyses of when (sub-)answers are determined in the CoT process, using synthetic arithmetic problems as a controlled testbed. Across a range of models and task difficulties, we found that models predominantly operated in a Talk-to-Think mode: they resolved the necessary subproblems during CoT generation phase. Moreover, causal intervention experiments revealed a strong recency bias linking hidden states to outputs, indicating that LLMs rely heavily on recent computations when generating explanations. This pattern further suggests that their internal reasoning flow largely aligns with the produced explanations.
Limitations
Comprehensiveness of experimental settings
Some experiments were conducted with a limited scope; for example, the experiments with various models in §˜4.3 are conducted only on the Level 3 task. Additionally, causal interventions (§˜5) are performed only with Qwen2.5-7B. Conducting our experiment with more models and tasks will further enhance the generalizability of the results.
Variety of task
We analyzed the internal reasoning patterns of language models using synthetic arithmetic reasoning tasks. The use of synthetic data allows for more detailed control compared to experiments on natural language tasks. However, vocabulary and expression diversity, for example, are limited compared to natural language tasks. Therefore, conducting similar analyses on reasoning tasks will verify whether the results of this study apply to other broader, realistic contexts as well. Additionally, in our study, we focus on a single reasoning-chain pattern, and it would be desirable to also conduct experiments using other reasoning chain strategies and formats. On the other hand, controlling the length and granularity of them is difficult because there are various options. Conducting experiments for other reasoning chain strategies and formats is expected to provide more general insights.
Probing methods
Interpreting internal mechanisms of LMs using probing have been actively conducted in our field Conneau et al. (2018); Tenney et al. (2019); Campbell et al. (2023); Li et al. (2023); however, there are criticisms regarding the validity of some probing approaches Liu et al. (2023); Burns et al. (2023). One way to overcome such concerns will be to analyze the generality of obtained results through more diverse methodologies Gurnee et al. (2023); Bricken et al. (2023).
Ethics statement
This paper will not raise particular ethical concerns, considering that (i) no human experiments were conducted, and (ii) our tasks do not involve ethically sensitive topics.
Appendix A Supplemental results
Model Level Task
Qwen2.5 (7B) 1 100
2 100
3 100
4 100
5 100
Qwen2.5 (14B) 3 100
Qwen2.5 (32B) 3 100
Qwen2.5-Math (7B) 3 100
Yi1.5 (9B) 3 100
Yi1.5 (34B) 3 100
Llama3.1 (8B) 3 100
Llama3.2 (3B) 3 97.6
Mistral-Nemo (12B) 3 99.6
Table 4: The performance of language models on the arithmetic reasoning tasks. The $\mathrm{Task}$ column shows the accuracy for the evaluation set (exact match).
A.1 Performance of language models in the arithmetic tasks
Table 4 shows the accuracy of language models on arithmetic reasoning tasks for each experimental setting. We computed the accuracy based on exact matches between the output, including the chain ($\hat{z}\oplus\hat{y}$), and the gold labels ($z\oplus y$). The accuracy for all models is nearly 100%, indicating that they are capable of solving the arithmetic reasoning tasks used in this experiment.
A.2 All probing results
Figures 9 through 53 present the probing results for all models and tasks discussed in this paper. Tables 6 through 20 summarize these results for thresholds ($\tau$) ranging from 0.85 to 0.95.
From these results, we observe trends similar to those described in §˜4.3 across many settings. However, for the smaller model Llama3.2 (3B), increasing the threshold $\tau$ often leads to cases where the accuracy does not reach the threshold (N/A). Nonetheless, a consistent pattern remains: $\mathrm{Acc}_{\mathrm{\prec CoT}}(v_{i})$ is low whereas $\mathrm{Acc}_{\mathrm{\succ CoT}}(v_{i})$ is high, indicating a Talk-to-Think mode.
Appendix B Hyperparameters
Table 5 shows the hyperparameters used for training the probes.
Train instances 10,000
Optimizer SGD Robbins (1951)
Learning rate $1.0\times 10^{-3}$ (constant)
Batch size 10,000
Epochs 10,000
Table 5: Hyperparameters for training the probe
Figure 6: Success and Unchanged rates for each grid when the final answer $y$ $({\underline{{\color[rgb]{.5,.5,.5}\mathrm{A}=}\textbf{6}}}_{\hskip 1.42271pt\mathbf{5}}$) is the target token. The Success rate heatmap at the top is the same as Figure 5.
Figure 7: Success rate and Unchanged rate for each grid when intervention was performed with $z_{17}$ (${\underline{{\color[rgb]{.5,.5,.5}\mathrm{A}=1+}5}}_{\hskip 1.42271pt\mathbf{4}}$) as the target token.
Figure 8: Success rate and Unchanged rate for each grid when intervention was performed with $z_{32}$ (${\underline{{\color[rgb]{.5,.5,.5}\mathrm{B}=}5}}_{\hskip 1.42271pt\mathbf{2}}$) as the target token.
Appendix C Addtional causal intervention results
Figures 7 and 8 show the causal intervention results for the target tokens $z_{17}$ and $z_{32}$, respectively. Here, in addition to the Success rate, we also present the Unchanged rate as a metric. The Unchanged rate indicates how frequently (%) the intervened output $\hat{\tilde{y}}$ remains the same as $y$. If this value is small, it indicates that the patched hidden states do not affect the output.
Appendix D Computational resources
We used NVIDIA A100 GPUs (40GB and 80GB memory) and NVIDIA H100 GPUs to conduct this study.
Appendix E Usage of AI assistants
For writing this paper and the source code for the experiments, we use AI assistants (e.g., ChatGPT, GitHub Copilot). However, the use is limited to purposes such as code completion, translation, text editing, and table creation, and all content is solely based on the authors’ ideas.
When ($\downarrow$) $\mathrm{Acc}$ ($\uparrow$)
Model Variable $^{*}_{\mathrm{eq}} tt^{*}$ $\mathrm{\prec CoT}$ $\mathrm{\succ CoT}$
Qwen2.5 (7B) $v_{1}$ ($\mathrm{A}$) 4 27 35.8 100
Qwen Team (2024) $v_{2}$ ($\mathrm{B}$) -2 -5 100 100
Qwen2.5 (14B) $v_{1}$ ($\mathrm{A}$) 4 27 36.9 100
Qwen Team (2024) $v_{2}$ ($\mathrm{B}$) -2 -5 100 100
Qwen2.5 (32B) $v_{1}$ ($\mathrm{A}$) 4 28 30.5 100
Qwen Team (2024) $v_{2}$ ($\mathrm{B}$) -2 -5 100 100
Qwen2.5-Math (7B) $v_{1}$ ($\mathrm{A}$) 4 27 41.8 100
Yang et al. (2024a) $v_{2}$ ($\mathrm{B}$) -2 -5 100 100
Yi1.5 (9B) $v_{1}$ ($\mathrm{A}$) 4 32 28.1 100
Young et al. (2024) $v_{2}$ ($\mathrm{B}$) -2 -5 100 100
Yi1.5 (34B) $v_{1}$ ($\mathrm{A}$) 4 31 22.9 100
Young et al. (2024) $v_{2}$ ($\mathrm{B}$) -2 -5 100 100
Llama3.1 (8B) $v_{1}$ ($\mathrm{A}$) 4 27 20.6 100
Dubey et al. (2024) $v_{2}$ ($\mathrm{B}$) -2 -5 100 100
Llama3.2 (3B) $v_{1}$ ($\mathrm{A}$) 4 28 21.8 100.0
Dubey et al. (2024) $v_{2}$ ($\mathrm{B}$) -2 -5 100 100
Mistral-Nemo (12B) $v_{1}$ ($\mathrm{A}$) 4 27 18.0 100
Mistral AI Team (2024) $v_{2}$ ($\mathrm{B}$) -2 -5 100 100
Table 6: Results for various models on the task Level 1 ($\tau=0.85$).
When ($\downarrow$) $\mathrm{Acc}$ ($\uparrow$)
Model Variable $^{*}_{\mathrm{eq}} tt^{*}$ $\mathrm{\prec CoT}$ $\mathrm{\succ CoT}$
Qwen2.5 (7B) $v_{1}$ ($\mathrm{A}$) 2 16 49.2 100
Qwen Team (2024) $v_{2}$ ($\mathrm{B}$) 5 35 21.2 100
Qwen2.5 (14B) $v_{1}$ ($\mathrm{A}$) 2 15 48.8 100
Qwen Team (2024) $v_{2}$ ($\mathrm{B}$) 5 36 21.5 100
Qwen2.5 (32B) $v_{1}$ ($\mathrm{A}$) 2 15 66.4 100
Qwen Team (2024) $v_{2}$ ($\mathrm{B}$) 5 36 21.3 100
Qwen2.5-Math (7B) $v_{1}$ ($\mathrm{A}$) 2 15 53.7 100
Yang et al. (2024a) $v_{2}$ ($\mathrm{B}$) 5 35 22.1 100
Yi1.5 (9B) $v_{1}$ ($\mathrm{A}$) 2 18 40.2 100
Young et al. (2024) $v_{2}$ ($\mathrm{B}$) 5 41 17.8 100
Yi1.5 (34B) $v_{1}$ ($\mathrm{A}$) 2 18 35.6 100
Young et al. (2024) $v_{2}$ ($\mathrm{B}$) 5 41 18.3 100
Llama3.1 (8B) $v_{1}$ ($\mathrm{A}$) 2 15 31.9 100
Dubey et al. (2024) $v_{2}$ ($\mathrm{B}$) 5 35 17.8 100
Llama3.2 (3B) $v_{1}$ ($\mathrm{A}$) 2 16 36.2 99.9
Dubey et al. (2024) $v_{2}$ ($\mathrm{B}$) 5 36 17.8 99.9
Mistral-Nemo (12B) $v_{1}$ ($\mathrm{A}$) 2 16 30.8 100
Mistral AI Team (2024) $v_{2}$ ($\mathrm{B}$) 5 36 17.8 100
Table 7: Results for various models on the task Level 2 ($\tau=0.85$).
When ($\downarrow$) $\mathrm{Acc}$ ($\uparrow$)
Model Variable $^{*}_{\mathrm{eq}} tt^{*}$ $\mathrm{\prec CoT}$ $\mathrm{\succ CoT}$
Qwen2.5 (7B) $v_{1}$ ($\mathrm{A}$) 5 35 17.9 100
Qwen Team (2024) $v_{2}$ ($\mathrm{B}$) 2 16 50.5 100
Qwen2.5 (14B) $v_{1}$ ($\mathrm{A}$) 5 35 17.8 100
Qwen Team (2024) $v_{2}$ ($\mathrm{B}$) 2 15 50.5 100
Qwen2.5 (32B) $v_{1}$ ($\mathrm{A}$) 5 36 17.8 100
Qwen Team (2024) $v_{2}$ ($\mathrm{B}$) 2 15 67.4 100
Qwen2.5-Math (7B) $v_{1}$ ($\mathrm{A}$) 5 35 18.6 100
Yang et al. (2024a) $v_{2}$ ($\mathrm{B}$) 2 15 56.1 100
Yi1.5 (9B) $v_{1}$ ($\mathrm{A}$) 5 41 17.8 100
Young et al. (2024) $v_{2}$ ($\mathrm{B}$) 2 18 36.9 100
Yi1.5 (34B) $v_{1}$ ($\mathrm{A}$) 5 41 22.4 100
Young et al. (2024) $v_{2}$ ($\mathrm{B}$) 2 18 37.4 100
Llama3.1 (8B) $v_{1}$ ($\mathrm{A}$) 5 35 26.0 100
Dubey et al. (2024) $v_{2}$ ($\mathrm{B}$) 2 16 29.6 100
Llama3.2 (3B) $v_{1}$ ($\mathrm{A}$) 5 36 17.8 93.2
Dubey et al. (2024) $v_{2}$ ($\mathrm{B}$) 2 16 33.2 95.4
Mistral-Nemo (12B) $v_{1}$ ($\mathrm{A}$) 5 36 17.8 100
Mistral AI Team (2024) $v_{2}$ ($\mathrm{B}$) 2 16 28.9 100
Table 8: Results for various models on the task Level 3 ($\tau=0.85$).
When ($\downarrow$) $\mathrm{Acc}$ ($\uparrow$)
Model Variable $^{*}_{\mathrm{eq}} tt^{*}$ $\mathrm{\prec CoT}$ $\mathrm{\succ CoT}$
Qwen2.5 (7B) $v_{1}$ ($\mathrm{A}$) 4 29 30.4 100
Qwen Team (2024) $v_{2}$ ($\mathrm{B}$) 2 15 27.2 100
$v_{3}$ ($\mathrm{C}$) N/A N/A 18.5 17.6
Qwen2.5 (14B) $v_{1}$ ($\mathrm{A}$) 5 35 18.9 100
Qwen Team (2024) $v_{2}$ ($\mathrm{B}$) 2 15 44.3 100
$v_{3}$ ($\mathrm{C}$) N/A N/A 40.4 26.8
Qwen2.5 (32B) $v_{1}$ ($\mathrm{A}$) 5 36 17.4 100
Qwen Team (2024) $v_{2}$ ($\mathrm{B}$) 2 15 62.8 100
$v_{3}$ ($\mathrm{C}$) N/A N/A 64.4 32.6
Qwen2.5-Math (7B) $v_{1}$ ($\mathrm{A}$) 5 35 17.2 100
Yang et al. (2024a) $v_{2}$ ($\mathrm{B}$) 2 15 55.6 100
$v_{3}$ ($\mathrm{C}$) N/A N/A 47.8 29.4
Yi1.5 (9B) $v_{1}$ ($\mathrm{A}$) 5 41 17.8 100
Young et al. (2024) $v_{2}$ ($\mathrm{B}$) 2 18 43.5 100
$v_{3}$ ($\mathrm{C}$) N/A N/A 36.7 21.2
Yi1.5 (34B) $v_{1}$ ($\mathrm{A}$) 5 40 19.3 100
Young et al. (2024) $v_{2}$ ($\mathrm{B}$) 2 18 40.8 100
$v_{3}$ ($\mathrm{C}$) N/A N/A 27.9 26.2
Llama3.1 (8B) $v_{1}$ ($\mathrm{A}$) 4 29 30.4 100
Dubey et al. (2024) $v_{2}$ ($\mathrm{B}$) 2 15 27.2 100
$v_{3}$ ($\mathrm{C}$) N/A N/A 18.5 17.6
Llama3.2 (3B) $v_{1}$ ($\mathrm{A}$) 5 36 26.2 91.7
Dubey et al. (2024) $v_{2}$ ($\mathrm{B}$) 2 16 29.1 98.7
$v_{3}$ ($\mathrm{C}$) N/A N/A 18.3 17.3
Mistral-Nemo (12B) $v_{1}$ ($\mathrm{A}$) 5 36 17.2 100
Mistral AI Team (2024) $v_{2}$ ($\mathrm{B}$) 2 16 29.9 100
$v_{3}$ ($\mathrm{C}$) N/A N/A 22.0 19.8
Table 9: Results for various models on the task Level 4 ($\tau=0.85$).
When ($\downarrow$) $\mathrm{Acc}$ ($\uparrow$)
Model Variable $^{*}_{\mathrm{eq}} tt^{*}$ $\mathrm{\prec CoT}$ $\mathrm{\succ CoT}$
Qwen2.5 (7B) $v_{1}$ ($\mathrm{A}$) 9 62 18.1 100
Qwen Team (2024) $v_{2}$ ($\mathrm{B}$) 6 42 22.6 100
$v_{3}$ ($\mathrm{C}$) 3 23 50.6 100
Qwen2.5 (14B) $v_{1}$ ($\mathrm{A}$) 9 63 18.1 98.8
Qwen Team (2024) $v_{2}$ ($\mathrm{B}$) 6 42 18.7 98.9
$v_{3}$ ($\mathrm{C}$) 3 23 42.2 100
Qwen2.5 (32B) $v_{1}$ ($\mathrm{A}$) 9 63 18.7 100
Qwen Team (2024) $v_{2}$ ($\mathrm{B}$) 6 43 22.6 100
$v_{3}$ ($\mathrm{C}$) 3 23 62.4 100
Qwen2.5-Math (7B) $v_{1}$ ($\mathrm{A}$) 9 62 18.1 100
Yang et al. (2024a) $v_{2}$ ($\mathrm{B}$) 6 42 22.6 100
$v_{3}$ ($\mathrm{C}$) 3 22 54.5 100
Yi1.5 (34B) $v_{1}$ ($\mathrm{A}$) 9 71 18.1 100
Young et al. (2024) $v_{2}$ ($\mathrm{B}$) 6 49 22.6 100
$v_{3}$ ($\mathrm{C}$) 3 26 41.2 100
Llama3.1 (8B) $v_{1}$ ($\mathrm{A}$) 9 62 16.0 99.5
Dubey et al. (2024) $v_{2}$ ($\mathrm{B}$) 6 43 20.0 99.5
$v_{3}$ ($\mathrm{C}$) 3 23 30.6 99.8
Llama3.2 (3B) $v_{1}$ ($\mathrm{A}$) N/A N/A 14.2 43.7
Dubey et al. (2024) $v_{2}$ ($\mathrm{B}$) N/A N/A 26.3 47.4
$v_{3}$ ($\mathrm{C}$) N/A N/A 37.7 71.7
Mistral-Nemo (12B) $v_{1}$ ($\mathrm{A}$) 9 63 18.1 99.9
Mistral AI Team (2024) $v_{2}$ ($\mathrm{B}$) 6 43 16.3 99.9
$v_{3}$ ($\mathrm{C}$) 3 23 32.0 99.9
Table 10: Results for various models on the task Level 5 ($\tau=0.85$).
When ($\downarrow$) $\mathrm{Acc}$ ($\uparrow$)
Model Variable $^{*}_{\mathrm{eq}} tt^{*}$ $\mathrm{\prec CoT}$ $\mathrm{\succ CoT}$
Qwen2.5 (7B) $v_{1}$ ($\mathrm{A}$) 4 27 35.8 100
Qwen Team (2024) $v_{2}$ ($\mathrm{B}$) -2 -5 100 100
Qwen2.5 (14B) $v_{1}$ ($\mathrm{A}$) 4 27 36.9 100
Qwen Team (2024) $v_{2}$ ($\mathrm{B}$) -2 -5 100 100
Qwen2.5 (32B) $v_{1}$ ($\mathrm{A}$) 4 28 30.5 100
Qwen Team (2024) $v_{2}$ ($\mathrm{B}$) -2 -5 100 100
Qwen2.5-Math (7B) $v_{1}$ ($\mathrm{A}$) 4 27 41.8 100
Yang et al. (2024a) $v_{2}$ ($\mathrm{B}$) -2 -5 100 100
Yi1.5 (9B) $v_{1}$ ($\mathrm{A}$) 4 32 28.1 100
Young et al. (2024) $v_{2}$ ($\mathrm{B}$) -2 -5 100 100
Yi1.5 (34B) $v_{1}$ ($\mathrm{A}$) 4 32 22.9 100
Young et al. (2024) $v_{2}$ ($\mathrm{B}$) -2 -5 100 100
Llama3.1 (8B) $v_{1}$ ($\mathrm{A}$) 4 27 20.6 100
Dubey et al. (2024) $v_{2}$ ($\mathrm{B}$) -2 -5 100 100
Llama3.2 (3B) $v_{1}$ ($\mathrm{A}$) 4 28 21.8 100.0
Dubey et al. (2024) $v_{2}$ ($\mathrm{B}$) -2 -5 100 100
Mistral-Nemo (12B) $v_{1}$ ($\mathrm{A}$) 4 28 18.0 100
Mistral AI Team (2024) $v_{2}$ ($\mathrm{B}$) -2 -5 100 100
Table 11: Results for various models on the task Level 1 ($\tau=0.90$).
When ($\downarrow$) $\mathrm{Acc}$ ($\uparrow$)
Model Variable $^{*}_{\mathrm{eq}} tt^{*}$ $\mathrm{\prec CoT}$ $\mathrm{\succ CoT}$
Qwen2.5 (7B) $v_{1}$ ($\mathrm{A}$) 2 16 49.2 100
Qwen Team (2024) $v_{2}$ ($\mathrm{B}$) 5 35 21.2 100
Qwen2.5 (14B) $v_{1}$ ($\mathrm{A}$) 2 16 48.8 100
Qwen Team (2024) $v_{2}$ ($\mathrm{B}$) 5 36 21.5 100
Qwen2.5 (32B) $v_{1}$ ($\mathrm{A}$) 2 16 66.4 100
Qwen Team (2024) $v_{2}$ ($\mathrm{B}$) 5 36 21.3 100
Qwen2.5-Math (7B) $v_{1}$ ($\mathrm{A}$) 2 15 53.7 100
Yang et al. (2024a) $v_{2}$ ($\mathrm{B}$) 5 35 22.1 100
Yi1.5 (9B) $v_{1}$ ($\mathrm{A}$) 2 18 40.2 100
Young et al. (2024) $v_{2}$ ($\mathrm{B}$) 5 41 17.8 100
Yi1.5 (34B) $v_{1}$ ($\mathrm{A}$) 2 18 35.6 100
Young et al. (2024) $v_{2}$ ($\mathrm{B}$) 5 41 18.3 100
Llama3.1 (8B) $v_{1}$ ($\mathrm{A}$) 2 15 31.9 100
Dubey et al. (2024) $v_{2}$ ($\mathrm{B}$) 5 35 17.8 100
Llama3.2 (3B) $v_{1}$ ($\mathrm{A}$) 2 16 36.2 99.9
Dubey et al. (2024) $v_{2}$ ($\mathrm{B}$) 5 36 17.8 99.9
Mistral-Nemo (12B) $v_{1}$ ($\mathrm{A}$) 2 16 30.8 100
Mistral AI Team (2024) $v_{2}$ ($\mathrm{B}$) 5 36 17.8 100
Table 12: Results for various models on the task Level 2 ($\tau=0.90$).
When ($\downarrow$) $\mathrm{Acc}$ ($\uparrow$)
Variable $^{*}_{\mathrm{eq}} tt^{*}$ $\mathrm{\prec CoT}$ $\mathrm{\succ CoT}$
Qwen2.5 (7B) $v_{1}$ ($\mathrm{A}$) 5 36 17.9 100
Qwen Team (2024) $v_{2}$ ($\mathrm{B}$) 2 16 50.5 100
Qwen2.5 (14B) $v_{1}$ ($\mathrm{A}$) 5 35 17.8 100
Qwen Team (2024) $v_{2}$ ($\mathrm{B}$) 2 16 50.5 100
Qwen2.5 (32B) $v_{1}$ ($\mathrm{A}$) 5 36 17.8 100
Qwen Team (2024) $v_{2}$ ($\mathrm{B}$) 2 15 67.4 100
Qwen2.5-Math (7B) $v_{1}$ ($\mathrm{A}$) 5 35 18.6 100
Yang et al. (2024a) $v_{2}$ ($\mathrm{B}$) 2 15 56.1 100
Yi1.5 (9B) $v_{1}$ ($\mathrm{A}$) 5 41 17.8 100
Young et al. (2024) $v_{2}$ ($\mathrm{B}$) 2 18 36.9 100
Yi1.5 (34B) $v_{1}$ ($\mathrm{A}$) 5 41 22.4 100
Young et al. (2024) $v_{2}$ ($\mathrm{B}$) 2 18 37.4 100
Llama3.1 (8B) $v_{1}$ ($\mathrm{A}$) 5 35 26.0 100
Dubey et al. (2024) $v_{2}$ ($\mathrm{B}$) 2 16 29.6 100
Llama3.2 (3B) $v_{1}$ ($\mathrm{A}$) 5 36 17.8 93.2
Dubey et al. (2024) $v_{2}$ ($\mathrm{B}$) 2 17 33.2 95.4
Mistral-Nemo (12B) $v_{1}$ ($\mathrm{A}$) 5 36 17.8 100
Mistral AI Team (2024) $v_{2}$ ($\mathrm{B}$) 2 16 28.9 100
Table 13: Results for various models on the task Level 3 ($\tau=0.90$).
When ($\downarrow$) $\mathrm{Acc}$ ($\uparrow$)
Model Variable $^{*}_{\mathrm{eq}} tt^{*}$ $\mathrm{\prec CoT}$ $\mathrm{\succ CoT}$
Qwen2.5 (7B) $v_{1}$ ($\mathrm{A}$) 5 35 17.2 100
Qwen Team (2024) $v_{2}$ ($\mathrm{B}$) 2 16 47.7 100
$v_{3}$ ($\mathrm{C}$) N/A N/A 43.7 23.7
Qwen2.5 (14B) $v_{1}$ ($\mathrm{A}$) 5 36 18.9 100
Qwen Team (2024) $v_{2}$ ($\mathrm{B}$) 2 15 44.3 100
$v_{3}$ ($\mathrm{C}$) N/A N/A 40.4 26.8
Qwen2.5 (32B) $v_{1}$ ($\mathrm{A}$) 5 36 17.4 100
Qwen Team (2024) $v_{2}$ ($\mathrm{B}$) 2 15 62.8 100
$v_{3}$ ($\mathrm{C}$) N/A N/A 64.4 32.6
Qwen2.5-Math (7B) $v_{1}$ ($\mathrm{A}$) 5 35 17.2 100
Yang et al. (2024a) $v_{2}$ ($\mathrm{B}$) 2 15 55.6 100
$v_{3}$ ($\mathrm{C}$) N/A N/A 47.8 29.4
Yi1.5 (9B) $v_{1}$ ($\mathrm{A}$) 5 41 17.8 100
Young et al. (2024) $v_{2}$ ($\mathrm{B}$) 2 18 43.5 100
$v_{3}$ ($\mathrm{C}$) N/A N/A 36.7 21.2
Yi1.5 (34B) $v_{1}$ ($\mathrm{A}$) 5 40 19.3 100
Young et al. (2024) $v_{2}$ ($\mathrm{B}$) 2 18 40.8 100
$v_{3}$ ($\mathrm{C}$) N/A N/A 27.9 26.2
Llama3.1 (8B) $v_{1}$ ($\mathrm{A}$) 5 35 30.4 100
Dubey et al. (2024) $v_{2}$ ($\mathrm{B}$) 2 16 27.2 100
$v_{3}$ ($\mathrm{C}$) N/A N/A 18.5 17.6
Llama3.2 (3B) $v_{1}$ ($\mathrm{A}$) 5 37 26.2 91.7
Dubey et al. (2024) $v_{2}$ ($\mathrm{B}$) 2 17 29.1 98.7
$v_{3}$ ($\mathrm{C}$) N/A N/A 18.3 17.3
Mistral-Nemo (12B) $v_{1}$ ($\mathrm{A}$) 5 36 17.2 100
Mistral AI Team (2024) $v_{2}$ ($\mathrm{B}$) 2 16 29.9 100
$v_{3}$ ($\mathrm{C}$) N/A N/A 22.0 19.8
Table 14: Results for various models on the task Level 4 ($\tau=0.90$).
When ($\downarrow$) $\mathrm{Acc}$ ($\uparrow$)
Model Variable $^{*}_{\mathrm{eq}} tt^{*}$ $\mathrm{\prec CoT}$ $\mathrm{\succ CoT}$
Qwen2.5 (7B) $v_{1}$ ($\mathrm{A}$) 9 63 18.1 100
Qwen Team (2024) $v_{2}$ ($\mathrm{B}$) 6 42 22.6 100
$v_{3}$ ($\mathrm{C}$) 3 23 50.6 100
Qwen2.5 (14B) $v_{1}$ ($\mathrm{A}$) 9 63 18.1 98.8
Qwen Team (2024) $v_{2}$ ($\mathrm{B}$) 6 42 18.7 98.9
$v_{3}$ ($\mathrm{C}$) 3 23 42.2 100
Qwen2.5 (32B) $v_{1}$ ($\mathrm{A}$) 9 63 18.7 100
Qwen Team (2024) $v_{2}$ ($\mathrm{B}$) 6 43 22.6 100
$v_{3}$ ($\mathrm{C}$) 3 23 62.4 100
Qwen2.5-Math (7B) $v_{1}$ ($\mathrm{A}$) 9 62 18.1 100
Yang et al. (2024a) $v_{2}$ ($\mathrm{B}$) 6 42 22.6 100
$v_{3}$ ($\mathrm{C}$) 3 22 54.5 100
Yi1.5 (34B) $v_{1}$ ($\mathrm{A}$) 9 72 18.1 100
Young et al. (2024) $v_{2}$ ($\mathrm{B}$) 6 49 22.6 100
$v_{3}$ ($\mathrm{C}$) 3 26 41.2 100
Llama3.1 (8B) $v_{1}$ ($\mathrm{A}$) 9 62 16.0 99.5
Dubey et al. (2024) $v_{2}$ ($\mathrm{B}$) 6 43 20.0 99.5
$v_{3}$ ($\mathrm{C}$) 3 23 30.6 99.8
Llama3.2 (3B) $v_{1}$ ($\mathrm{A}$) N/A N/A 14.2 43.7
Dubey et al. (2024) $v_{2}$ ($\mathrm{B}$) N/A N/A 26.3 47.4
$v_{3}$ ($\mathrm{C}$) N/A N/A 37.7 71.7
Mistral-Nemo (12B) $v_{1}$ ($\mathrm{A}$) 9 63 18.1 99.9
Mistral AI Team (2024) $v_{2}$ ($\mathrm{B}$) 6 43 16.3 99.9
$v_{3}$ ($\mathrm{C}$) 3 23 32.0 99.9
Table 15: Results for various models on the task Level 5 ($\tau=0.90$).
When ($\downarrow$) $\mathrm{Acc}$ ($\uparrow$)
Model Variable $^{*}_{\mathrm{eq}} tt^{*}$ $\mathrm{\prec CoT}$ $\mathrm{\succ CoT}$
Qwen2.5 (7B) $v_{1}$ ($\mathrm{A}$) 4 27 35.8 100
Qwen Team (2024) $v_{2}$ ($\mathrm{B}$) -2 -5 100 100
Qwen2.5 (14B) $v_{1}$ ($\mathrm{A}$) 4 28 36.9 100
Qwen Team (2024) $v_{2}$ ($\mathrm{B}$) -2 -5 100 100
Qwen2.5 (32B) $v_{1}$ ($\mathrm{A}$) 4 28 30.5 100
Qwen Team (2024) $v_{2}$ ($\mathrm{B}$) -2 -5 100 100
Qwen2.5-Math (7B) $v_{1}$ ($\mathrm{A}$) 4 27 41.8 100
Yang et al. (2024a) $v_{2}$ ($\mathrm{B}$) -2 -5 100 100
Yi1.5 (9B) $v_{1}$ ($\mathrm{A}$) 4 32 28.1 100
Young et al. (2024) $v_{2}$ ($\mathrm{B}$) -2 -5 100 100
Yi1.5 (34B) $v_{1}$ ($\mathrm{A}$) 4 32 22.9 100
Young et al. (2024) $v_{2}$ ($\mathrm{B}$) -2 -5 100 100
Llama3.1 (8B) $v_{1}$ ($\mathrm{A}$) 4 27 20.6 100
Dubey et al. (2024) $v_{2}$ ($\mathrm{B}$) -2 -5 100 100
Llama3.2 (3B) $v_{1}$ ($\mathrm{A}$) 4 28 21.8 100.0
Dubey et al. (2024) $v_{2}$ ($\mathrm{B}$) -2 -5 100 100
Mistral-Nemo (12B) $v_{1}$ ($\mathrm{A}$) 4 28 17.9 100
Mistral AI Team (2024) $v_{2}$ ($\mathrm{B}$) -2 -5 100 100
Table 16: Results for various models on the task Level 1 ($\tau=0.95$).
When ($\downarrow$) $\mathrm{Acc}$ ($\uparrow$)
Model Variable $^{*}_{\mathrm{eq}} tt^{*}$ $\mathrm{\prec CoT}$ $\mathrm{\succ CoT}$
Qwen2.5 (7B) $v_{1}$ ($\mathrm{A}$) 2 16 49.2 100
Qwen Team (2024) $v_{2}$ ($\mathrm{B}$) 5 36 21.2 100
Qwen2.5 (14B) $v_{1}$ ($\mathrm{A}$) 2 16 48.8 100
Qwen Team (2024) $v_{2}$ ($\mathrm{B}$) 5 36 21.5 100
Qwen2.5 (32B) $v_{1}$ ($\mathrm{A}$) 2 16 66.4 100
Qwen Team (2024) $v_{2}$ ($\mathrm{B}$) 5 36 21.3 100
Qwen2.5-Math (7B) $v_{1}$ ($\mathrm{A}$) 2 15 53.7 100
Yang et al. (2024a) $v_{2}$ ($\mathrm{B}$) 5 36 22.1 100
Yi1.5 (9B) $v_{1}$ ($\mathrm{A}$) 2 18 40.2 100
Young et al. (2024) $v_{2}$ ($\mathrm{B}$) 5 41 17.8 100
Yi1.5 (34B) $v_{1}$ ($\mathrm{A}$) 2 18 35.6 100
Young et al. (2024) $v_{2}$ ($\mathrm{B}$) 5 41 18.3 100
Llama3.1 (8B) $v_{1}$ ($\mathrm{A}$) 2 15 31.9 100
Dubey et al. (2024) $v_{2}$ ($\mathrm{B}$) 5 35 17.8 100
Llama3.2 (3B) $v_{1}$ ($\mathrm{A}$) 2 16 36.2 99.9
Dubey et al. (2024) $v_{2}$ ($\mathrm{B}$) 5 36 17.8 99.9
Mistral-Nemo (12B) $v_{1}$ ($\mathrm{A}$) 2 16 30.8 100
Mistral AI Team (2024) $v_{2}$ ($\mathrm{B}$) 5 36 17.8 100
Table 17: Results for various models on the task Level 2 ($\tau=0.95$).
When ($\downarrow$) $\mathrm{Acc}$ ($\uparrow$)
Model Variable $^{*}_{\mathrm{eq}} tt^{*}$ $\mathrm{\prec CoT}$ $\mathrm{\succ CoT}$
Qwen2.5 (7B) $v_{1}$ ($\mathrm{A}$) 5 36 17.9 100
Qwen Team (2024) $v_{2}$ ($\mathrm{B}$) 2 16 50.5 100
Qwen2.5 (14B) $v_{1}$ ($\mathrm{A}$) 5 36 17.8 100
Qwen Team (2024) $v_{2}$ ($\mathrm{B}$) 2 16 50.5 100
Qwen2.5 (32B) $v_{1}$ ($\mathrm{A}$) 5 36 17.8 100
Qwen Team (2024) $v_{2}$ ($\mathrm{B}$) 2 15 67.4 100
Qwen2.5-Math (7B) $v_{1}$ ($\mathrm{A}$) 5 35 18.6 100
Yang et al. (2024a) $v_{2}$ ($\mathrm{B}$) 2 15 56.1 100
Yi1.5 (9B) $v_{1}$ ($\mathrm{A}$) 5 41 17.8 100
Young et al. (2024) $v_{2}$ ($\mathrm{B}$) 2 18 36.9 100
Yi1.5 (34B) $v_{1}$ ($\mathrm{A}$) 5 41 22.4 100
Young et al. (2024) $v_{2}$ ($\mathrm{B}$) 2 18 37.4 100
Llama3.1 (8B) $v_{1}$ ($\mathrm{A}$) 5 35 26.0 100
Dubey et al. (2024) $v_{2}$ ($\mathrm{B}$) 2 16 29.6 100
Llama3.2 (3B) $v_{1}$ ($\mathrm{A}$) N/A N/A 17.8 93.2
Dubey et al. (2024) $v_{2}$ ($\mathrm{B}$) 2 17 33.2 95.4
Mistral-Nemo (12B) $v_{1}$ ($\mathrm{A}$) 5 36 17.8 100
Mistral AI Team (2024) $v_{2}$ ($\mathrm{B}$) 2 16 28.9 100
Table 18: Results for various models on the task Level 3 ($\tau=0.95$).
When ($\downarrow$) $\mathrm{Acc}$ ($\uparrow$)
Model Variable $^{*}_{\mathrm{eq}} tt^{*}$ $\mathrm{\prec CoT}$ $\mathrm{\succ CoT}$
Qwen2.5 (7B) $v_{1}$ ($\mathrm{A}$) 5 36 17.2 100
Qwen Team (2024) $v_{2}$ ($\mathrm{B}$) 2 16 47.7 100
$v_{3}$ ($\mathrm{C}$) N/A N/A 43.7 23.7
Qwen2.5 (14B) $v_{1}$ ($\mathrm{A}$) 5 36 18.9 100
Qwen Team (2024) $v_{2}$ ($\mathrm{B}$) 2 15 44.3 100
$v_{3}$ ($\mathrm{C}$) N/A N/A 40.4 26.8
Qwen2.5 (32B) $v_{1}$ ($\mathrm{A}$) 5 36 17.4 100
Qwen Team (2024) $v_{2}$ ($\mathrm{B}$) 2 16 62.8 100
$v_{3}$ ($\mathrm{C}$) N/A N/A 64.4 32.6
Qwen2.5-Math (7B) $v_{1}$ ($\mathrm{A}$) 5 35 17.2 100
Yang et al. (2024a) $v_{2}$ ($\mathrm{B}$) 2 15 55.6 100
$v_{3}$ ($\mathrm{C}$) N/A N/A 47.8 29.4
Yi1.5 (9B) $v_{1}$ ($\mathrm{A}$) 5 41 17.8 100
Young et al. (2024) $v_{2}$ ($\mathrm{B}$) 2 18 43.5 100
$v_{3}$ ($\mathrm{C}$) N/A N/A 36.7 21.2
Yi1.5 (34B) $v_{1}$ ($\mathrm{A}$) 5 41 19.3 100
Young et al. (2024) $v_{2}$ ($\mathrm{B}$) 2 18 40.8 100
$v_{3}$ ($\mathrm{C}$) N/A N/A 27.9 26.2
Llama3.1 (8B) $v_{1}$ ($\mathrm{A}$) 5 35 30.4 100
Dubey et al. (2024) $v_{2}$ ($\mathrm{B}$) 2 16 27.2 100
$v_{3}$ ($\mathrm{C}$) N/A N/A 18.5 17.6
Llama3.2 (3B) $v_{1}$ ($\mathrm{A}$) N/A N/A 26.2 91.7
Dubey et al. (2024) $v_{2}$ ($\mathrm{B}$) 2 17 29.1 98.7
$v_{3}$ ($\mathrm{C}$) N/A N/A 18.3 17.3
Mistral-Nemo (12B) $v_{1}$ ($\mathrm{A}$) 5 36 17.2 100
Mistral AI Team (2024) $v_{2}$ ($\mathrm{B}$) 2 16 29.9 100
$v_{3}$ ($\mathrm{C}$) N/A N/A 22.0 19.8
Table 19: Results for various models on the task Level 4 ($\tau=0.95$).
When ($\downarrow$) $\mathrm{Acc}$ ($\uparrow$)
Model Variable $^{*}_{\mathrm{eq}} tt^{*}$ $\mathrm{\prec CoT}$ $\mathrm{\succ CoT}$
Qwen2.5 (7B) $v_{1}$ ($\mathrm{A}$) 9 63 18.1 100
Qwen Team (2024) $v_{2}$ ($\mathrm{B}$) 6 43 22.6 100
$v_{3}$ ($\mathrm{C}$) 3 23 50.6 100
Qwen2.5 (14B) $v_{1}$ ($\mathrm{A}$) 9 63 18.1 98.8
Qwen Team (2024) $v_{2}$ ($\mathrm{B}$) 6 43 18.7 98.9
$v_{3}$ ($\mathrm{C}$) 3 23 42.2 100
Qwen2.5 (32B) $v_{1}$ ($\mathrm{A}$) 9 63 18.7 100
Qwen Team (2024) $v_{2}$ ($\mathrm{B}$) 6 43 22.6 100
$v_{3}$ ($\mathrm{C}$) 3 23 62.4 100
Qwen2.5-Math (7B) $v_{1}$ ($\mathrm{A}$) 9 63 18.1 100
Yang et al. (2024a) $v_{2}$ ($\mathrm{B}$) 6 42 22.6 100
$v_{3}$ ($\mathrm{C}$) 3 22 54.5 100
Yi1.5 (34B) $v_{1}$ ($\mathrm{A}$) 9 72 18.1 100
Young et al. (2024) $v_{2}$ ($\mathrm{B}$) 6 49 22.6 100
$v_{3}$ ($\mathrm{C}$) 3 26 41.2 100
Llama3.1 (8B) $v_{1}$ ($\mathrm{A}$) 9 62 16.0 99.5
Dubey et al. (2024) $v_{2}$ ($\mathrm{B}$) 6 43 20.0 99.5
$v_{3}$ ($\mathrm{C}$) 3 23 30.6 99.8
Llama3.2 (3B) $v_{1}$ ($\mathrm{A}$) N/A N/A 14.1 43.7
Dubey et al. (2024) $v_{2}$ ($\mathrm{B}$) N/A N/A 26.3 47.4
$v_{3}$ ($\mathrm{C}$) N/A N/A 37.7 71.7
Mistral-Nemo (12B) $v_{1}$ ($\mathrm{A}$) 9 63 18.1 99.9
Mistral AI Team (2024) $v_{2}$ ($\mathrm{B}$) 6 43 16.3 99.9
$v_{3}$ ($\mathrm{C}$) 3 23 32.0 99.9
Table 20: Results for various models on the task Level 5 ($\tau=0.95$).
Figure 9: Probing results when Qwen2.5-7B solves Level 1.
Figure 10: Probing results when Qwen2.5-7B solves Level 2.
Figure 11: Probing results when Qwen2.5-7B solves Level 3.
Figure 12: Probing results when Qwen2.5-7B solves Level 4.
Figure 13: Probing results when Qwen2.5-7B solves Level 5.
Figure 14: Probing results when Qwen2.5-14B solves Level 1.
Figure 15: Probing results when Qwen2.5-14B solves Level 2.
Figure 16: Probing results when Qwen2.5-14B solves Level 3.
Figure 17: Probing results when Qwen2.5-14B solves Level 4.
Figure 18: Probing results when Qwen2.5-14B solves Level 5.
Figure 19: Probing results when Qwen2.5-32B solves Level 1.
Figure 20: Probing results when Qwen2.5-32B solves Level 2.
Figure 21: Probing results when Qwen2.5-32B solves Level 3.
Figure 22: Probing results when Qwen2.5-32B solves Level 4.
Figure 23: Probing results when Qwen2.5-32B solves Level 5.
Figure 24: Probing results when Qwen2.5-Math-7B solves Level 1.
Figure 25: Probing results when Qwen2.5-Math-7B solves Level 2.
Figure 26: Probing results when Qwen2.5-Math-7B solves Level 3.
Figure 27: Probing results when Qwen2.5-Math-7B solves Level 4.
Figure 28: Probing results when Qwen2.5-Math-7B solves Level 5.
Figure 29: Probing results when Yi-1.5-9B solves Level 1.
Figure 30: Probing results when Yi-1.5-9B solves Level 2.
Figure 31: Probing results when Yi-1.5-9B solves Level 3.
Figure 32: Probing results when Yi-1.5-9B solves Level 4.
Figure 33: Probing results when Yi-1.5-9B solves Level 5.
Figure 34: Probing results when Yi-1.5-34B solves Level 1.
Figure 35: Probing results when Yi-1.5-34B solves Level 2.
Figure 36: Probing results when Yi-1.5-34B solves Level 3.
Figure 37: Probing results when Yi-1.5-34B solves Level 4.
Figure 38: Probing results when Yi-1.5-34B solves Level 5.
Figure 39: Probing results when Llama-3.1-8B solves Level 1.
Figure 40: Probing results when Llama-3.1-8B solves Level 2.
Figure 41: Probing results when Llama-3.1-8B solves Level 3.
Figure 42: Probing results when Llama-3.1-8B solves Level 4.
Figure 43: Probing results when Llama-3.1-8B solves Level 5.
Figure 44: Probing results when Llama-3.2-3B solves Level 1.
Figure 45: Probing results when Llama-3.2-3B solves Level 2.
Figure 46: Probing results when Llama-3.2-3B solves Level 3.
Figure 47: Probing results when Llama-3.2-3B solves Level 4.
Figure 48: Probing results when Llama-3.2-3B solves Level 5.
Figure 49: Probing results when Mistral-Nemo-Base-2407 solves Level 1.
Figure 50: Probing results when Mistral-Nemo-Base-2407 solves Level 2.
Figure 51: Probing results when Mistral-Nemo-Base-2407 solves Level 3.
Figure 52: Probing results when Mistral-Nemo-Base-2407 solves Level 4.
Figure 53: Probing results when Mistral-Nemo-Base-2407 solves Level 5.